OpenStudy (anonymous):

eliminate the parameter t. find a rectangular equation for the plane cure defined by the parametric equations. x = 6 cos t y = 6 sin t 0<= t <=2n

OpenStudy (anonymous):

do you know what \[(\cos(t), sin(t))\] looks like?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

they are points on the unit circle you have points on the circle with radius 6

OpenStudy (anonymous):

okay I just didn't know if it would be x^2 - y^2 = 6 or x^2 + y^2 = 6

OpenStudy (anonymous):

it is neither

OpenStudy (anonymous):

oh so you have to square the 6 which would be 36 right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so then would it be add or subtract

OpenStudy (anonymous):

if you square and add you get \[x^2+y^2=36\cos^2(t)+36\sin^2(t)=36(\cos^2(t)+\sin^2(t))=36\] so it is \(x^2+y^2=36\)

OpenStudy (anonymous):

thank you